OpenStudy (anonymous):

Identify the solution of the inequality |x + 3| - 3 >/ 1

hero (hero):

I'm assuming you mean: \(|x + 3| - 3 \ge 1\)

OpenStudy (anonymous):

Yes.

hero (hero):

Hint: Add 3 to both sides

OpenStudy (anonymous):

\[x + 3 \ge 4 \]

hero (hero):

No, you can't get rid of the absolute value sign yet: \(|x+3| \ge 4\)

OpenStudy (anonymous):

How do you get rid of it then?

hero (hero):

Hint: \(|x + a| \ge c\) becomes \( -c < x + a < c \)

OpenStudy (anonymous):

\[-4 <x+3<4\] Now do you add it all up and set everything equal to X?

hero (hero):

Actually, let me correct my formula

hero (hero):

\(|x + a| \ge c \) becomes \( -c \ge x + a \ge c\)

hero (hero):

Sorry, I'm used to writing these a certain way and also I make typing errors sometimes. But that what you see above is the correct formula.

OpenStudy (anonymous):

\[-7 \le x \le1\] Would that be the correct solution?

hero (hero):

For one, the inequality symbols are wrong. That's the reason why I corrected myself.

hero (hero):

Double check the formula again.

hero (hero):

When you do it the way I showed you, you will end up with \(-7 \ge x \ge 1\)

hero (hero):

But at this point, you're probably realizing that this cannot be true.

OpenStudy (anonymous):

\[x \le -7 or x \ge 1\]

hero (hero):

May I ask how you figured that out.

hero (hero):

I was going to explain that what you posted is the correct answer.

OpenStudy (anonymous):

Isn't it because it would look like \[-7 \le x \ge1\]

hero (hero):

No, and you still haven't explained how you figured it out. You would never write it like that.

hero (hero):

But that might give me an idea of how to write these in the future.

hero (hero):

I'm almost certain you cheated somehow.

hero (hero):

I was supposed to explain to you how to arrive at that answer.

hero (hero):

There's no way you could have figured that out on your own.

OpenStudy (anonymous):

I remember some of what I learned in school, not all of it but certain parts a little. I'm trying to figure out how I remembered it.

hero (hero):

Well, anyway, thanks to you, I can re-write my general formula.

hero (hero):

\(| x + a | \ge c \) becomes \(-c \le x + a \ge c\)

hero (hero):

A little unconventional, but it works

hero (hero):

Actually, I just realized that the general formula may not work in every case.

hero (hero):

I'll have to test it out.